OpenStudy (anonymous):

Find the y-intercept of the line whose equation is 3x - 4y = -2.

OpenStudy (anonymous):

to find the y intercept, sub in 0 for x 3x - 4y = -2 3(0) - 4y = -2 -4y = -2 y = -2/-4 y = 1/2

OpenStudy (anonymous):

thank you so much!(:

OpenStudy (anonymous):

no problem...just remember, to find y intercept, sub in 0 for x. To find the x intercept, sub in 0 for y